Two firms, DeSoto Cab (D) and Uber (U), sell taxi cab rides in San Francisco. Suppose the demand curve for DeSoto is Dp = 60 - PD + 0.5pu while the demand curve for Uber is given by Du = 60 - Pu + 0.5pd. Suppose DeSoto has marginal cost of $25 and Uber has a marginal cost of $35. The two firms compete by choosing price (i.e. Bertrand competition).

(a) Derive each firm's best response (or reaction) functions.

(b) If Uber sets a price of $20, what price will DeSoto set and will it make any profits?

(c) Draw the demand curve for DeSoto if Uber sets a price of $20. Include its marginal revenue curve, marginal cost "curve," and the optimal price from part (b).

(d) What is the Bertrand-Nash equilibrium?
Answer #1

Answer Desoto Cab: Dp = 60-10 +0.5Pu Uber Cab : Du = 60-Pu +0.500 MC, 125 MC = $35 (2) Now - profit (x) = TR-TC an [P-Mc D ToBeol Response function: Desoto Can : P = 85 +0.5 Pu Uber cab: Pu - 95 +0.5 PRO (b) of Pu- $20, then Pp - 85 +0.5x20 -85 +5 Po(0) TRO= Pox Dp (TR: Total Revenue) Dp = 60-80 +0.5Pu Op = 60-Po + 0.5X20 = 70 -Po Po=70-Dp TRO = Dp (To-Dp Marginal Revenue2. - MC - $25 0 2 2 ho t e Suantito. Dp to (d) Bertrand - Nosh (quelibolum: Prices where Best funcion interses each othen Bro& Pu= 95 +0.5 (58) » Pu = 47.5+ 14.5 - [Pu = 62] Therefore, Bertrand -Nash Equilibrium is Po= 58, Pu = 62]